solve the following equation :2(2x+1)+5=3(2x+1)

Found 2 solutions by lynnlo, lenny460:
Answer by lynnlo(4176)
Answer by lenny460(1073)

solve the following equation :2(2x+1)+5=3(2x+1)

2 * 2x = 4x
2 * 1 = 2
3 * 2x = 6x
3 * 1 = 3
4x + 2 + 5 = 6x + 3
Combine like terms:
2 + 5 = 7
Therefore:
4x + 7 = 6x + 3
Subtract 6x from each side of the equation.
4x + 7 - 6x = 6x + 3 - 6x
Combine like terms:
4x - 6x = -2x
6x - 6x = 0
-2x + 7 = 3
Subtract 7 from each side of the equation.
-2x + 7 - 7 = 3 - 7
Combine like terms:
7 - 7 = 0
3 - 7 = -4
Therefore:
-2x = -4
Divide each side by -2
-2x/-2 = -4/-2
The Answer:
x = 2



Lennox Obuong
Algebra Student
Email: obuong3@aol.com